You want to have $3 million in real dollars in an account when you retire in 40 years. The nominal return on your investment is 13 percent and the inflation rate is 3.5 percent. What real amount must you deposit each year to achieve your goal?

  • $8,463.06

  • $2,959.44

  • $8,886.21

  • $8,801.58

  • $8,039.91
Answer #1

Real rate=(1+nominal rate)/(1+inflation)-1=1.13/1.035-1=9.17874%

Real amount=Future Real Value*Real rate/((1+real rate)^n-1)=3000000*9.17874%/(1.0917874^40-1)=8463.06